when the second number is subtracted from twice the first number the result is 3 the sum of three times the first number and twice the second number is 8 what are the numbers
Answer by richwmiller(17219)

a is the first and b is the second
when the second number is subtracted from twice the first number the result is 3

2a-b=3
the sum of three times the first number and twice the second number is 8 what are the numbers
3a+2b=8

2a-b=3
3a+2b=8
multiply first by 2
4a-2b=6
3a+2b=8
Add them.
7a=14
a=2
now sub into original first
2*2-b=3
4-b=3
b=1
2 and 1 are the numbers